### v1

# On the complex moment problem as a dynamic inverse problem for a discrete system

Abstract. We consider the complex moment problem, that is the problem of constructing a positive Borel measure on $\mathbb{C}$ from a given set of moments. We relate this problem to the dynamic inverse problem for the discrete system associated with the complex Jacobi matrix. We show how the characterization of dynamic inverse data in solving the inverse problem provides sufficient conditions for solving the complex moment problem.
1. Introduction.
The complex moment problem is: given a set of complex numbers $s_{0},s_{1},s_{2},\ldots$, to find a Borel measure $d\rho$ on $\mathbb{C}$ such that
(1.1) $s_{k}=\int_{\mathbb{C}}\lambda^{k}\,d\rho(\lambda),\quad k=0,1,2,\ldots.$
If such a measure exists $s_{0},s_{1},s_{2},\ldots$ are called moments of this measure. The truncated moment problem is the problem of finding a measure that satisfies a finite number of moment equalities (1.1) for $k=0,1,\ldots,2N-2$ for some $N\in\mathbb{N}$.
For a given sequence of complex numbers $\{a_{1},a_{2},\ldots\}$, $\{b_{1},b_{2},\ldots\}$, $a_{i}\not=0$, we define the complex Jacobi matrix:
(1.2) $A=\begin{pmatrix}b_{1}&a_{1}&0&0&0&\ldots\\ a_{1}&b_{2}&a_{2}&0&0&\ldots\\ 0&a_{2}&b_{3}&a_{3}&0&\ldots\\ \ldots&\ldots&\ldots&\ldots&\ldots&\ldots\end{pmatrix}$.
For $N\in\mathbb{N}$, by $A^{N}$ we denote the $N\times N$ Jacobi matrix which is a block of (1.2) consisting of the intersection of first $N$ columns with first $N$ rows of $A$.
Associated with the matrix $A$ and additional parameter $\mathbb{C}\ni a_{0}\not=0$ is a dynamical system with discrete time:
(1.3) $\begin{cases}u_{n,t+1}+u_{n,t-1}-a_{n}u_{n+1,t}-a_{n-1}u_{n-1,t}-b_{n}u_{n,t}=0,\quad n,t\in\mathbb{N},\\ u_{n,-1}=u_{n,0}=0,\quad n\in\mathbb{N},\\ u_{0,t}=f_{t},\quad t\in\mathbb{N}\cup\{0\},\end{cases}$
which is a discrete analogue of the dynamical systems governed by a wave equation on the semi-axis [1, 3]. By analogy with the continuous problems [2], we consider the complex sequence $f=(f_{0},f_{1},\ldots)$ as a boundary control. The solution to (1.3) we denote by $u^{f}_{n,t}$. We also consider the dynamical system associated with finite matrix $A_{N}$:
(1.4) $\begin{cases}v_{n,t+1}+v_{n,t-1}-a_{n}v_{n+1,t}-a_{n-1}v_{n-1,t}-b_{n}v_{n,t}=0,\,\,t\in\mathbb{N}_{0},\,\,n\in 1,\ldots,N,\\ v_{n,\,-1}=v_{n,\,0}=0,\quad n=1,2,\ldots,N+1,\\ v_{0,\,t}=f_{t},\quad v_{N+1,\,t}=0,\quad t\in\mathbb{N}_{0},\end{cases}$
which is a natural analog of the dynamical systems governed by a wave equation on the interval, the solution to (1.4) is denoted by $v^{f}$.
Fixing $T\in\mathbb{N}$, we associate the response operator with (1.3), which maps the control $f=(f_{0},\ldots f_{T-1})$ to $u^{f}_{1,t}$:
(1.5) $\left(R^{T}f\right)_{t}:=u^{f}_{1,t},\quad t=1,\ldots,T.$
In the second section using the Autonne-Takagi [11] factorization, we derive a special ”spectral representation” for the solution of (1.4).
In the third section we present results on the dynamic inverse problem (IP) for the system (1.3) in accordance with [8]. This problems is a natural discrete analogue of the IP for the wave equation on the half-axis, where the dynamic Dirichlet to Neumann map is used as inverse data, see [2]. The IP for the dynamical system (1.3) with real Jacobi matrix is considered in [5, 6]. The connections between the IP for the system with real Jacobi matrix and classical moment problems are described in [7, 9]. The IP for the dynamical system with complex Jacobi matrix is considered in [8].
In last section we describe connections between the IPs for (1.4) and (1.3) and complex moment problem and introduce a discrete Borel measure on $\mathbb{C}$, concentrated on the finite set of point, associated with (1.4), which is a solution to the truncated complex moment problem. Letting $N$ tend to infinity in (1.4) we obtain a sequence of measures that converges to a solution to the complex moment problem.
Another approach to the complex moment problem was proposed in [12], where the author used the generalized spectral function introduced in [4] as a basic tool and obtained sufficient conditions on the moment sequence under the condition that it is bounded. Note that our method do not require the boundedness of $s_{k}$, $k\geqslant 0$.
If $H\in C^{n\times n}$ then by $H^{*}$ we denote the matrix conjugate to $H$ and recall that $H^{*}=\overline{H^{\top}}$, where $H^{\top}$ is a transpose of $H$.
2. Discrete dynamical system. Forward problem, special representation of the solution.
We first derive a special representation (Fourier-type expansion) of the the solution to (1.4). Note that another representation (Duhamel-type), important for solving the IP obtained in [8], is used in the next section. Our special representation is based on the following Autonne-Takagi [11] factorization:
Theorem 1.
Let $H\in\mathbb{C}^{n\times n}$ be a complex symmetric matrix: $H^{*}=\overline{H}$, then there exists a unitary matrix $U$ such that
(2.1) $UHU^{\top}=D=\begin{pmatrix}d_{1}&0&\ldots&0\\ 0&d_{2}&\ldots&0\\ \ldots&\ldots&\ldots&\ldots\\ 0&0&\ldots&d_{n}\end{pmatrix}$,
where $d_{i}\geqslant 0,$ $i=1\ldots,n$.
We apply this theorem for the matrix $A^{N}$: there exist unitary $U^{N}$ such that $U^{N}A^{N}\left(U^{N}\right)^{\top}$ as in the theorem. Note that in our case all $d_{i}>0$, since otherwise $A^{N}u=0$ implies $\det A^{N}=0$, which in turn implies linear dependence of some rows or columns of $A^{N}$, which is impossible. At the same time some of the $d_{i}$ can coincide. In the latter case we modify the unitary matrix as follows: assume for example that $d_{1}=d_{2}$, then, taking
$U_{1}^{N}:=\begin{pmatrix}1&0&\ldots&0\\ 0&e^{i\frac{\varphi_{1}}{2}}&\ldots&0\\ \ldots&\ldots&\ldots&\ldots\\ 0&0&\ldots&1\end{pmatrix}U^{N},$
we obtain that
$U_{1}^{N}A^{N}\left(U_{1}^{N}\right)^{\top}=\begin{pmatrix}d_{1}&0&\ldots&0\\ 0&d_{2}e^{i\varphi}&\ldots&0\\ \ldots&\ldots&\ldots&\ldots\\ 0&0&\ldots&d_{n}\end{pmatrix}$.
Using these arguments we can make all $d_{i}$ in the representation (2.1) distinct (but some of them can be complex). To summarize what we did so far, in what follows we choose unitary $U$ (we drop $N$) such that
(2.2) $UA^{N}U^{\top}=\begin{pmatrix}d_{1}&0&\ldots&0\\ 0&d_{2}&\ldots&0\\ \ldots&\ldots&\ldots&\ldots\\ 0&0&\ldots&d_{n}\end{pmatrix},\quad d_{i}\in\mathbb{C}\backslash\{0\},d_{i}\not=d_{j},\,i\not=j,\,i,j=1\ldots,N.$
We introduce the vectors
$e_{i}=\begin{pmatrix}0\\ \cdot\\ 1\\ \cdot\\ 0\end{pmatrix},\quad\text{with $1$ on i-th place},i=1,\ldots,N.$
Then
$UA^{N}U^{\top}e_{i}=d_{i}e_{i},$
we multiply the equality above from the left by $U^{*}$ and get
$A^{N}U^{\top}e_{i}=d_{i}U^{*}e_{i}=d_{i}\overline{U^{\top}}e_{i},$
Introducing the notation
$U^{\top}=\left(\hat{u}^{1}\,|\,\hat{u}^{2}\,|\,\ldots\,|\,\hat{u}^{N}\right),\quad\hat{u}^{i}=\begin{pmatrix}\hat{u}^{i}_{1}\\ {\hat{u}}^{i}_{2}\\ \ldots\\ {\hat{u}}^{i}_{n}\end{pmatrix}$
that is $\hat{u}^{i}$ is a column in the matrix $U^{\top}$, we see that $\hat{u}^{i}$ satisfies:
(2.3) $A^{N}\hat{u}^{i}e_{i}=d_{i}\overline{\hat{u}^{i}},\quad A^{N}\begin{pmatrix}\hat{u}^{i}_{1}\\ {\hat{u}}^{i}_{2}\\ \ldots\\ {\hat{u}}^{i}_{n}\end{pmatrix}=d_{i}\begin{pmatrix}\overline{\hat{u}^{i}_{1}}\\ \overline{{u}^{i}_{2}}\\ \ldots\\ \overline{\hat{u}^{i}_{n}}\end{pmatrix}$.
Note that the first components of all vectors are non-zero:
$\hat{u}^{i}_{1}\not=0,\quad i=1,\ldots,N,$
otherwise it immediately follows from (2.3) that $\hat{u}^{i}=0$.
Thus, we can introduce the vectors which we use in the Fourier-type expansion for the solution to (1.4) in the following way:
$u^{i}=\begin{pmatrix}0\\ \frac{\hat{u}^{i}_{1}}{\hat{u}^{i}_{1}}\\ \ldots\\ \frac{\hat{u}^{i}_{N}}{\hat{u}^{i}_{1}}\\ 0\end{pmatrix},\quad i=1,\ldots,N,$
so we formally add two values: $u^{i}_{0}=u^{i}_{N+1}=0$ and normalize vectors such that $u^{i}_{1}=1$.
We take some vector $(y_{0},y_{1},\ldots y_{N},y_{N+1})$, multiply the equation in (1.4) by $y_{n}$, sum up and change the order of summation:
$0=\sum_{n=1}^{N}\left(v_{n,t+1}y_{n}+v_{n,t-1}y_{n}-a_{n-1}v_{n,t}y_{n-1}-a_{n}v_{n,t}y_{n+1}-b_{n}v_{n,t}y_{n}\right)$
$-a_{N}v_{N+1,t}y_{N}-a_{0}v_{0,t}y_{1}+a_{0}v_{1,t}y_{0}+a_{N}v_{N,t}y_{N+1}.$
Now we take $\left(y_{0},y_{1},\ldots y_{N},y_{N+1}\right)=\left(u^{i}_{0},u^{i}_{1},\ldots u^{i}_{N},u^{i}_{N+1}\right)$ and evaluate counting $u^{i}_{1}=1$ and the values at $n=0$ and $n=N+1$:
$0=\sum_{n=1}^{N}\left(v_{n,t+1}u^{i}_{n}+v_{n,t-1}u^{i}_{n}-v_{n,t}\left(a_{n-1}y^{i}_{n-1}+a_{n}vu^{i}_{n+1}+b_{n}u^{i}_{n}\right)\right)-a_{0}v_{0,t}u^{i}_{1}.$
Counting (2.3) we get:
(2.4) $0=\sum_{n=1}^{N}\left(v_{n,t+1}u^{i}_{n}+v_{n,t-1}u^{i}_{n}-d_{i}v_{n,t}\overline{u^{i}_{n}}\right)-a_{0}f_{t}\quad i=1,\ldots,N.$
Now we look for the solution to (1.4) in the form:
(2.5) $v_{n,t}=\begin{cases}\sum_{k=1}^{N}c^{k}_{t}\overline{u^{k}_{n}},\\ f_{t},\quad n=0.\end{cases}$
Plugging (2.5) into (2.4) we have:
(2.6) $\sum_{n=1}^{N}\left(\sum_{k=1}^{N}\left(c_{t+1}^{k}\overline{u^{k}_{n}}+c_{t-1}^{k}\overline{u^{k}_{n}}\right)u_{n}^{i}-d_{i}\sum_{k=1}^{N}c^{k}_{t}\overline{u^{k}_{n}}\overline{u^{i}_{n}}\right)=a_{0}f_{t}.$
Introducing the notations
(2.7) $\sum_{n=1}^{N}\overline{u^{k}_{n}}u_{n}^{i}=\delta_{ki}\rho_{i},\quad i=1,\ldots,N,$
(2.8) $H_{ki}=\sum_{n=1}^{N}\overline{u^{k}_{n}}\overline{u_{n}^{i}},\quad k,i=1,\ldots,N,$
we rewrite (2.6) as:
$\left(c_{t+1}^{k}+c_{t-1}^{k}\right)\delta_{ki}\rho_{i}-d_{i}\sum_{k=1}^{N}c^{k}_{t}H_{ki}=a_{0}f_{t},\quad t,k,i=1\ldots,N.$
So we see that $c^{i}_{t}$ are determined from
(2.9) $c_{t+1}^{i}+c_{t-1}^{i}-\frac{d_{i}}{\rho_{i}}\sum_{k=1}^{N}c^{k}_{t}H_{ki}=\frac{a_{0}}{\rho_{i}}f_{t},\quad t,i=1,\ldots,N.$
We look for the solution to (2.9) in the form:
(2.10) $c_{t}^{i}=\frac{a_{0}}{\rho_{i}}\sum_{l=0}^{t}f_{l}T_{t-l}^{(i)},\quad t,i=1,\ldots,N.$
Plugging this representation into (2.9) we get:
$\frac{a_{0}}{\rho_{i}}\left(\sum_{l=0}^{t}f_{l}T_{t+1-l}^{(i)}+\sum_{l=0}^{t-1}f_{l}T_{t-1-l}^{(i)}-\sum_{k=1}^{N}d_{i}\frac{H_{ki}}{\rho_{k}}\sum_{l=0}^{t}f_{l}T_{t-l}^{(i)}\right)=\frac{a_{0}}{\rho_{i}}f_{t},$
changing the order of summation (at this point we use the additional value $T_{-1}^{(i)}$) we come to
(2.11) $\sum_{l=0}^{t}f_{l}\left(T_{t+1-l}^{(i)}+T_{t-1-l}^{(i)}-\left(\sum_{k=1}^{N}d_{i}\frac{H_{ki}}{\rho_{k}}\right)T_{t-l}^{(i)}\right)+f_{t+1}T_{0}^{(i)}-f_{t}T_{-1}^{(i)}=f_{t}.$
We introduce the notation
(2.12) $\omega_{i}=\sum_{k=1}^{N}d_{i}\frac{H_{ki}}{\rho_{k}},\quad i=1,\ldots,N,$
then (2.11) holds if $T^{(i)}_{t}$ satisfies
(2.13) $\begin{cases}T_{t+1}^{(i)}+T_{t-1}^{(i)}-\omega_{i}T_{t}^{(i)}=0,\quad t=0,1\ldots,N,\\ T_{0}^{(i)}=0,\,\,T_{-1}^{(i)}=-1.\end{cases}$
Or, in other words, $T_{t}^{(i)}$ are simply the Chebyshev polynomials of the first kind evaluated at points $\omega_{i}$: $T^{(i)}_{t}=T_{t}(\omega_{i})$.
3. Inverse problem for discrete dynamical system associated with complex Jacobi matrices.
In this section we outline the results of the IP for (1.3) according to [8].
We fix some positive integer $T$ and denote by $\mathcal{F}^{T}$ the outer space of the system (1.3), the space of controls: $\mathcal{F}^{T}:=\mathbb{C}^{T}$, $f\in\mathcal{F}^{T}$, $f=(f_{0},\ldots,f_{T-1})$, $f,g\in\mathcal{F}^{T}$, $(f,g)_{\mathcal{F}^{T}}=\sum_{k=0}^{T-1}f_{k}\overline{g_{k}}$. And let $\mathcal{F}^{\infty}=\left\{(f_{0},f_{1},\ldots)\,|\,f_{i}\in\mathbb{C},\,i=0,1,\ldots\right\}$, so $\mathcal{F}^{\infty}$ is the set of complex sequences. The following representation formula for the solution to (1.3) can be considered as an analogue of a Duhamel representation formula for the initial-boundary value problem for the wave equation with a potential on the half-line [1].
Lemma 1.
A solution to (1.3) admits the representation
(3.1) $u^{f}_{n,t}=\prod_{k=0}^{n-1}a_{k}f_{t-n}+\sum_{s=n}^{t-1}w_{n,s}f_{t-s-1},\quad n,t\in\mathbb{N},$
where $w_{n,s}$ satisfies the Goursat problem
$\left\{\begin{array}[]{l}w_{n,s+1}+w_{n,s-1}-a_{n}w_{n+1,s}-a_{n-1}w_{n-1,s}-b_{n}w_{n,s}=\\ =-\delta_{s,n}(1-a_{n}^{2})\prod_{k=0}^{n-1}a_{k},\,n,s\in\mathbb{N},\,\,s>n,\\ w_{n,n}-b_{n}\prod_{k=0}^{n-1}a_{k}-a_{n-1}w_{n-1,n-1}=0,\quad n\in\mathbb{N},\\ w_{0,t}=0,\quad t\in\mathbb{N}_{0}.\end{array}\right$.
Definition 1.
For $f,g\in\mathcal{F}^{\infty}$ we define the convolution $c=f*g\in\mathcal{F}^{\infty}$ by the formula
$c_{t}=\sum_{s=0}^{t}f_{s}g_{t-s},\quad t\in\mathbb{N}\cup\{0\}.$
Let us introduce an analog of the dynamic response operator (dynamic Dirichlet-to-Neumann map) [2] for the system (1.3):
Definition 2.
The response operator $R^{T}:\mathcal{F}^{T}\mapsto\mathbb{C}^{T}$ for the system (1.3) is defined by (1.5)
The response vector is the convolution kernel of the response operator, $r=(r_{0},r_{1},\ldots,r_{T-1})=(a_{0},w_{1,1},w_{1,2},\ldots w_{1,T-1})$, according to (3.1):
(3.2) $\left(R^{T}f\right)_{t}=u^{f}_{1,t}=a_{0}f_{t-1}+\sum_{s=1}^{t-1}w_{1,s}f_{t-1-s}\quad t=1,\ldots,T.$
$\left(R^{T}f\right)=r*f_{\cdot-1}.$
By choosing the special control $f=\delta=(1,0,0,\ldots)$, the kernel of the response operator can be determined as
(3.3) $\left(R^{T}\delta\right)_{t}=u^{\delta}_{1,t}=r_{t-1},\quad t=1,2,\ldots.$
The inverse problem considered in [8] consists in recovering the Jacobi matrix (i.e. the sequences $\{a_{1},a_{2},\ldots\}$, $\{b_{1},b_{2},\ldots\}$) and $a_{0}$ from the response operator.
In what follows we use the same notations for operators and for matrices of these operators.
We introduce the inner space of dynamical system (1.3) $\mathcal{H}^{T}:=\mathbb{C}^{T}$, $h\in\mathcal{H}^{T}$, $h=(h_{1},\ldots,h_{T})$ with the inner product $h,l\in\mathcal{H}^{T}$, $(h,g)_{\mathcal{H}^{T}}=\sum_{k=1}^{T}h_{k}\overline{g_{k}}$. The control operator $W^{T}:\mathcal{F}^{T}\mapsto\mathcal{H}^{T}$ is defined by the rule
(3.4) $W^{T}f:=u^{f}_{n,T},\quad n=1,\ldots,T.$
From (3.1) we deduce the representation for $W^{T}$:
$\left(W^{T}f\right)_{n}=u^{f}_{n,T}=\prod_{k=0}^{n-1}a_{k}f_{T-n}+\sum_{s=n}^{T-1}w_{n,s}f_{T-s-1},\quad n=1,\ldots,T.$
The following statement is equivalent to boundary controllability of (1.3).
Lemma 2.
The operator $W^{T}$ is an isomorphism between $\mathcal{F}^{T}$ and $\mathcal{H}^{T}$.
Along with the system (1.3) we consider an auxiliary system associated with the complex conjugate matrix $\overline{A}$:
(3.5) $\begin{cases}v_{n,t+1}+v_{n,t-1}-\overline{a_{n}}v_{n+1,t}-\overline{a_{n-1}}v_{n-1,t}-\overline{b_{n}}v_{n,t}=0,\quad n,t\in\mathbb{N},\\ v_{n,-1}=v_{n,0}=0,\quad n\in\mathbb{N},\\ v_{0,t}=f_{t},\quad t\in\mathbb{N}\cup\{0\}.\end{cases}$
The objects corresponding to the system (3.5) are marked with the symbol #\#. Direct calculations show:
Lemma 3.
The control and response operators of the system #\# a related with control and response operators of the original system by the relations
(3.6) $W^{T}_{\#}=\overline{W^{T}},\quad R^{T}_{\#}=\overline{R^{T}},$
that is, the matrix of $W^{T}_{\#}$ and the response vector $r_{\#}$ are complex conjugate of the matrix of $W^{T}$ and the vector $r$.
For systems (1.3), (3.5) we introduce the connecting operator $C^{T}:\mathcal{F}^{T}\mapsto\mathcal{F}^{T}$ via the bilinear form: for arbitrary $f,g\in\mathcal{F}^{T}$ we define
(3.7) $\left(C^{T}f,g\right)_{\mathcal{F}^{T}}=\left(u^{f}_{\cdot,T},v^{g}_{\cdot,T}\right)_{\mathcal{H}^{T}}=\left(W^{T}f,W^{T}_{\#}g\right)_{\mathcal{H}^{T}}.$
The following statement is crucial for solving the dynamic inverse problem:
Theorem 2.
The connecting operator $C^{T}$ is an isomorphism in $\mathcal{F}^{T}$, it admits the representation in terms of inverse data:
(3.8) $C^{T}=a_{0}C^{T}_{ij},\quad C^{T}_{ij}=\sum_{k=0}^{T-\max{i,j}}r_{|i-j|+2k},\quad r_{0}=a_{0},$
$C^{T}=\begin{pmatrix}r_{0}+r_{2}+\ldots+r_{2T-2}&r_{1}+r_{3}+\ldots+r_{2T-3}&\cdot&r_{T}+r_{T-2}&r_{T-1}\\ r_{1}+r_{3}+\ldots+r_{2T-3}&r_{0}+r_{2}+\ldots+r_{2T-4}&\cdot&\ldots&r_{T-2}\\ \cdot&\cdot&\cdot&\cdot&\cdot\\ r_{T-3}+r_{T-1}+r_{T+1}&\ldots&\cdot&r_{1}+r_{3}&r_{2}\\ r_{T}+r_{T-2}&\ldots&\cdot&r_{0}+r_{2}&r_{1}\\ r_{T-1}&r_{T-2}&\cdot&r_{1}&r_{0}\end{pmatrix}$.
The relations (3.6) imply the following
Remark 1.
The connecting operator is complex symmetric:
$\left(C^{T}\right)^{*}=\overline{C^{T}},\quad\text{or}\quad\left(C^{T}\right)^{\top}=C^{T}.$
3.1. Inverse problem.
Due to the finite speed of wave propagation in (1.3) the solution $u^{f}$ depends on the coefficients $a_{n},b_{n}$ as follows:
Remark 2.
For $M\in\mathbb{N}$, $u^{f}_{M-1,M}$ depends on $\{a_{0},\ldots,a_{M-1}\}$, $\{b_{1},\ldots,b_{M-1}\}$, the response $R^{2T}$ (or, what is equivalent, the response vector $(r_{0},r_{1},\ldots,r_{2T-2})$) depends on $\{a_{0},\ldots,a_{T-1}\}$, $\{b_{1},\ldots,b_{T}\}$.
Thus the natural set up of the dynamic IP for (1.3): by the given operator $R^{2T}$ to recover $\{a_{0},\ldots,a_{T-1}\}$ and $\{b_{1},\ldots,b_{T-1}\}$.
We also note that $a_{0}=r_{0}$, which follows from (3.2).
In [8] the authors proposed two methods for recovering the coefficients $(a_{k})^{2}$, $b_{k},$ $k=1,\ldots$. To recover $a_{k}$ it is necessary to use additional information, such as the sequence of signs. Note that the results obtained for dynamic inverse data in [8] corresponds to results obtained for spectral inverse data in [4].
Note that the impossibility of recovering $a_{k}$ is not the weak point of the method, but a feature of the problem:
Theorem 3.
  • 1)
    For $f\in\mathcal{F}$ the value $u^{f}_{n,t}$ is odd with respect to $a_{1},a_{2},\ldots a_{n-1}$ and even with respect to $a_{n},a_{n+1},\ldots$.
  • 2)
    The response vector depends on $\left(a_{1}\right)^{2},\left(a_{2}\right)^{2},\ldots$
Now we set up a question: can one determine whether a vector $(r_{0},r_{1},r_{2},\ldots,r_{2T-2})$ is a response vector for dynamical system (1.3) with some $(a_{0},\ldots,a_{T-1})$ $(b_{1},\ldots,b_{T-1})$? The answer is the following theorem.
Theorem 4.
The vector $(r_{0},r_{1},r_{2},\ldots,r_{2T-2})$ is a response vector for the dynamical system (1.3) if and only if the complex symmetric matrix $C^{T-k}$, $k=0,1,\ldots,T-1$ constructed by (3.8) is an isomorphism in $\mathcal{F}^{T-k}$.
4. Complex moment problem and dynamic inverse problem.
In what follows we will assume that additional parameter $a_{0}=1$.
With the dynamical system (1.4) one can also associate the control, response and connecting operators $W^{T}_{N},$ $R^{T}_{N}$ and $C^{T}_{N}$ with the same formulas (3.4), (1.5), (3.7), where instead of $u^{f}$ one should use $v^{f}$.
Remark 2 in particular implies that
(4.1) $R^{2N-2}=R^{2N-2}_{N},$
$u^{f}_{n,\,t}=v^{f}_{n,\,t},\quad n\leqslant t\leqslant N,$
$W^{T}=W^{T}_{N},\quad C^{T}=C^{T}_{N},\quad T\leqslant N.$
By choosing the special control $f=\delta=(1,0,0,\ldots)$, the kernel of a response operator can be determined as (cf. (3.3)):
$\left(R^{T}_{N}\delta\right)_{t}=v^{\delta}_{1,t}=r^{N}_{t-1},\quad t=1,2,\ldots.$
Note that from (4.1) it follows that
$r_{t}=r_{t}^{N},\quad t=0,1,\ldots,2N-2.$
Thus for special control $f=\delta$, using (2.5), (2.10), (2.13) one have that:
(4.2) $r_{t-1}^{N}=v^{\delta}_{1,t}=\sum_{k=1}^{N}c^{k}_{t}\overline{u^{k}_{1}}=\sum_{k=1}^{N}c^{k}_{t}=\sum_{k=1}^{N}\frac{1}{\rho_{k}}T_{t}(\omega_{k}),\quad t=1,2,\ldots,$
where $\rho_{k}$ and $\omega_{k}$ are defined in (2.7) and (2.12).
We introduce a discrete measure $d\rho^{N}$ on $\mathbb{C}$, concentrated on the set of points $\left\{\omega_{k}\right\}_{k=1}^{N}$, by definition we set
(4.3) $d\rho^{N}(\{\omega_{k}\})=\frac{1}{\rho_{k}},$
so that at points $\omega_{k}$ it has weights $\frac{1}{\rho_{k}}$. Then we can rewrite (4.2) in a form which resembles the spectral representation of dynamic inverse data (see [6, 7, 9]):
Proposition 1.
The dynamic response vector of the system (1.4) admits the following representation:
(4.4) $r_{t-1}^{N}=\int_{\mathbb{C}}T_{t}(\lambda)\,d\rho^{N}(\lambda),\quad t=1,2,\ldots.$
With a set of moments (1.1) we associate the following Hankel matrices:
(4.5) $S^{n}:=\begin{pmatrix}s_{2n-2}&s_{2n-3}&\ldots&s_{n-1}\\ s_{2n-3}&\ldots&\ldots&\ldots\\ \cdot&\cdot&\ldots&s_{1}\\ s_{n-1}&\ldots&s_{1}&s_{0}\end{pmatrix},\quad n=2,3,\ldots.$
We also introduce the matrix $J_{n}\in\mathbb{C}^{n\times n}$
$J_{n}=\begin{pmatrix}0&0&0&\ldots&1\\ 0&0&0&\ldots&0\\ \cdot&\cdot&\cdot&\cdot&\cdot\\ 0&\ldots&1&0&0\\ \cdot&\cdot&\cdot&\cdot&\cdot\\ 1&0&0&0&0\end{pmatrix},\quad n=2,3,\ldots.$
In [7, 9] the authors obtained the following
Proposition 2.
The elements of the response vector (4.4) of the system (1.4) are related to the moments $s_{k}=\int_{\mathbb{C}}\lambda^{k}\,d\rho^{N}(\lambda)$ by the following rule:
(4.6) $\begin{pmatrix}r_{0}^{N}\\ r_{1}^{N}\\ \ldots\\ r_{n-1}^{N}\end{pmatrix}=\Lambda_{n}\begin{pmatrix}s_{0}\\ s_{1}\\ \ldots\\ s_{n-1}\end{pmatrix},\quad n=1,2,\ldots,2N-1.$
where the entries of the matrix $\Lambda_{n}\in\mathbb{R}^{n\times n}$ are given by
$\{\Lambda_{n}\}_{ij}=a_{ij}=\begin{cases}0,\quad\text{if $i<j$},\\ 0,\quad\text{if $i+j$ is odd,}\\ E_{\frac{i+j}{2}}^{j}(-1)^{\frac{i+j}{2}+j},\quad\text{otherwise}\end{cases}$
where $E_{n}^{k}$ are binomial coefficients.
The following relation holds:
$C^{N}=\widetilde{\Lambda}_{N}S^{N}\left(\widetilde{\Lambda}_{N}\right)^{*},\quad\widetilde{\Lambda}_{N}:=J_{N}\Lambda_{N}J_{N}.$
Thus, we arrived at the following procedure for solving the truncated complex moment problem, i.e. the problem of finding a measure that satisfies a finite number of moment equalities (1.1) for $s_{0},s_{1},\ldots s_{2N-2}\in\mathbb{C}$:
Procedure 1.
The solution to the truncated moment problem can be constructed by performing the following steps
  • 1)
    Go from $s_{0},s_{1},\ldots s_{2N-2}$ to $r_{0},\ldots,r_{2N-2}$ using the formula (4.6).
  • 2)
    Check if $r_{0},r_{1},\ldots r_{2N-2}$ satisfy the conditions of Theorem 4, then the elements of the matrix $A^{N}$ can be calculated using the formulas from [8].
  • 3)
    Calculate parameters $d_{i}$, $\rho_{i}$, $H_{ki}$, $\omega_{i}$ using factorization (2.2) and formulas (2.7), (2.8), (2.12).
  • 4)
    Construct a measure that solves the truncated moment problem by (4.3)
The possibility of constructing a solution to the truncated complex moment problem, described above, can be used as a basis for the following
Theorem 5.
If the moments $s_{0},s_{1},\ldots$ are such that the Hankel matrix $S^{T}$ (4.5) is non-singular for every $T\in N$, there exist a solution to the complex moment problem (1.1).
Proof.
Taking an arbitrary $T\in\mathbb{N}$, and using the fact that for $T^{\prime}\leqslant T$ the corresponding $C^{T^{\prime}}$ and $S^{T^{\prime}}$ are simultaneously non-singular, we use Procedure 1 and obtain the measure $d\rho^{T}$ that solves the truncated moment problem for the set of moments $s_{0},s_{1},\ldots,s_{2T-2}$.
Then for arbitrary polynomial $P\in C[\lambda]$, $P(\lambda)=\sum_{k=0}^{M}a_{k}\lambda^{k}$, we have that
(4.7) $\int_{\mathbb{C}}P(\lambda)\,d\rho^{N}(\lambda)\longrightarrow_{N\to\infty}\sum_{k=0}^{M}a_{k}s_{k}.$
this follows from the fact that by the construction we have that
$\int_{\mathbb{C}}P(\lambda)\,d\rho^{N}(\lambda)=\sum_{k=0}^{M}a_{k}s_{k},\quad\text{if}\quad 2N-2>M.$
Convergence (4.7) means that $d\rho^{N}$ converges *-weakly as $N\to\infty$ to some measure $d\rho$, and this $d\rho$ is a solution to the complex moment problem. (1.1).
Note that our method allows one to associate a certain measure with the Jacobi matrix (1.2). The study of this measure, its connection with the generalized spectral function [4] as well as related functions [10] will be the subject of the forthcoming publications.

### v2

@@
 We relate this problem to the dynamic inverse problem for the discrete system associated with the complex Jacobi matrix.
-We show how the characterization of dynamic inverse data in solving the inverse problem provides sufficient conditions for solving the complex moment problem.
+We propose a method that allows one to construct a discrete measure which is a solution to the truncated moment problem, we also show how the characterization of dynamic inverse data in solving the inverse problem provides sufficient conditions for solving the full complex moment problem.
 1. Introduction.
@@
 In last section we describe connections between the IPs for (1.4) and (1.3) and complex moment problem and introduce a discrete Borel measure on $\mathbb{C}$, concentrated on the finite set of point, associated with (1.4), which is a solution to the truncated complex moment problem.
-Letting $N$ tend to infinity in (1.4) we obtain a sequence of measures that converges to a solution to the complex moment problem.
+We emphasize that we propose a method that allows one to construct a discrete measure, i.e.
+find points of its support and determine the masses at these points.
+Letting $N$ tend to infinity in (1.4) we obtain a sequence of measures that converges to a solution to the full complex moment problem.
 Another approach to the complex moment problem was proposed in [12], where the author used the generalized spectral function introduced in [4] as a basic tool and obtained sufficient conditions on the moment sequence under the condition that it is bounded.
@@
 Let $H\in\mathbb{C}^{n\times n}$ be a complex symmetric matrix: $H^{*}=\overline{H}$, then there exists a unitary matrix $U$ such that
-(2.1) $UHU^{\top}=D=\begin{pmatrix}d_{1}&0&\ldots&0\\ 0&d_{2}&\ldots&0\\ \ldots&\ldots&\ldots&\ldots\\ 0&0&\ldots&d_{n}\end{pmatrix}$,
-where $d_{i}\geqslant 0,$ $i=1\ldots,n$.
+(2.1) $UHU^{\top}=D=\begin{pmatrix}\hat{d}_{1}&0&\ldots&0\\ 0&\hat{d}_{2}&\ldots&0\\ \ldots&\ldots&\ldots&\ldots\\ 0&0&\ldots&\hat{d}_{n}\end{pmatrix}$,
+where $\hat{d}_{i}\geqslant 0,$ $i=1\ldots,n$.
 We apply this theorem for the matrix $A^{N}$: there exist unitary $U^{N}$ such that $U^{N}A^{N}\left(U^{N}\right)^{\top}$ as in the theorem.
-Note that in our case all $d_{i}>0$, since otherwise $A^{N}u=0$ implies $\det A^{N}=0$, which in turn implies linear dependence of some rows or columns of $A^{N}$, which is impossible.
-At the same time some of the $d_{i}$ can coincide.
-In the latter case we modify the unitary matrix as follows: assume for example that $d_{1}=d_{2}$, then, taking
+Note that in our case all $\hat{d}_{i}>0$, since otherwise $A^{N}u=0$ implies $\det A^{N}=0$, which in turn implies linear dependence of some rows or columns of $A^{N}$, which is impossible.
+At the same time some of the $\hat{d}_{i}$ can coincide.
+In the latter case we modify the unitary matrix as follows: assume for example that $\hat{d}_{1}=\hat{d}_{2}$, then, taking
 $U_{1}^{N}:=\begin{pmatrix}1&0&\ldots&0\\ 0&e^{i\frac{\varphi_{1}}{2}}&\ldots&0\\ \ldots&\ldots&\ldots&\ldots\\ 0&0&\ldots&1\end{pmatrix}U^{N},$
 we obtain that
-$U_{1}^{N}A^{N}\left(U_{1}^{N}\right)^{\top}=\begin{pmatrix}d_{1}&0&\ldots&0\\ 0&d_{2}e^{i\varphi}&\ldots&0\\ \ldots&\ldots&\ldots&\ldots\\ 0&0&\ldots&d_{n}\end{pmatrix}$.
-Using these arguments we can make all $d_{i}$ in the representation (2.1) distinct (but some of them can be complex).
+$U_{1}^{N}A^{N}\left(U_{1}^{N}\right)^{\top}=\begin{pmatrix}\hat{d}_{1}&0&\ldots&0\\ 0&\hat{d}_{2}e^{i\varphi}&\ldots&0\\ \ldots&\ldots&\ldots&\ldots\\ 0&0&\ldots&\hat{d}_{n}\end{pmatrix}$.
+Using these arguments we can make all $\hat{d}_{i}$ in the representation (2.1) distinct (but some of them can be complex).
 To summarize what we did so far, in what follows we choose unitary $U$ (we drop $N$) such that
-(2.2) $UA^{N}U^{\top}=\begin{pmatrix}d_{1}&0&\ldots&0\\ 0&d_{2}&\ldots&0\\ \ldots&\ldots&\ldots&\ldots\\ 0&0&\ldots&d_{n}\end{pmatrix},\quad d_{i}\in\mathbb{C}\backslash\{0\},d_{i}\not=d_{j},\,i\not=j,\,i,j=1\ldots,N.$
+(2.2) $UA^{N}U^{\top}=\begin{pmatrix}\hat{d}_{1}&0&\ldots&0\\ 0&\hat{d}_{2}&\ldots&0\\ \ldots&\ldots&\ldots&\ldots\\ 0&0&\ldots&\hat{d}_{n}\end{pmatrix},\quad\hat{d}_{i}\in\mathbb{C}\backslash\{0\},\hat{d}_{i}\not=\hat{d}_{j},\,i\not=j,\,i,j=1\ldots,N.$
 We introduce the vectors
@@
 Then
-$UA^{N}U^{\top}e_{i}=d_{i}e_{i},$
+$UA^{N}U^{\top}e_{i}=\hat{d}_{i}e_{i},$
 we multiply the equality above from the left by $U^{*}$ and get
-$A^{N}U^{\top}e_{i}=d_{i}U^{*}e_{i}=d_{i}\overline{U^{\top}}e_{i},$
+$A^{N}U^{\top}e_{i}=\hat{d}_{i}U^{*}e_{i}=\hat{d}_{i}\overline{U^{\top}}e_{i},$
 Introducing the notation
@@
 that is $\hat{u}^{i}$ is a column in the matrix $U^{\top}$, we see that $\hat{u}^{i}$ satisfies:
-(2.3) $A^{N}\hat{u}^{i}e_{i}=d_{i}\overline{\hat{u}^{i}},\quad A^{N}\begin{pmatrix}\hat{u}^{i}_{1}\\ {\hat{u}}^{i}_{2}\\ \ldots\\ {\hat{u}}^{i}_{n}\end{pmatrix}=d_{i}\begin{pmatrix}\overline{\hat{u}^{i}_{1}}\\ \overline{{u}^{i}_{2}}\\ \ldots\\ \overline{\hat{u}^{i}_{n}}\end{pmatrix}$.
+(2.3) $A^{N}\hat{u}^{i}=\hat{d}_{i}\overline{\hat{u}^{i}},\quad A^{N}\begin{pmatrix}\hat{u}^{i}_{1}\\ {\hat{u}}^{i}_{2}\\ \ldots\\ {\hat{u}}^{i}_{n}\end{pmatrix}=\hat{d}_{i}\begin{pmatrix}\overline{\hat{u}^{i}_{1}}\\ \overline{{u}^{i}_{2}}\\ \ldots\\ \overline{\hat{u}^{i}_{n}}\end{pmatrix}$.
 Note that the first components of all vectors are non-zero:
@@
 $u^{i}=\begin{pmatrix}0\\ \frac{\hat{u}^{i}_{1}}{\hat{u}^{i}_{1}}\\ \ldots\\ \frac{\hat{u}^{i}_{N}}{\hat{u}^{i}_{1}}\\ 0\end{pmatrix},\quad i=1,\ldots,N,$
-so we formally add two values: $u^{i}_{0}=u^{i}_{N+1}=0$ and normalize vectors such that $u^{i}_{1}=1$.
+so we formally add two values: $u^{i}_{0}=u^{i}_{N+1}=0$ and normalize vectors such that $u^{i}_{1}=1$. In this case we have (see (2.3))
+$A^{N}\begin{pmatrix}\frac{\hat{u}^{i}_{1}}{\hat{u}^{i}_{1}}\\ \frac{\hat{u}^{i}_{2}}{\hat{u}^{i}_{1}}\\ \ldots\\ \frac{\hat{u}^{i}_{n}}{\hat{u}^{i}_{1}}\end{pmatrix}={d}_{i}\begin{pmatrix}\frac{\overline{\hat{u}^{i}_{1}}}{\overline{\hat{u}^{i}_{1}}}\\ \frac{\overline{\hat{u}^{i}_{2}}}{\overline{\hat{u}^{i}_{1}}}\\ \ldots\\ \frac{\overline{\hat{u}^{i}_{n}}}{{\overline{\hat{u}^{i}_{1}}}}\end{pmatrix}$,
+where we introduced the notation
+(2.4) $d_{i}:=\hat{d_{i}}\frac{\overline{\hat{u}^{i}_{1}}}{{\hat{u}^{i}_{1}}},\quad i=1,\ldots,N.$
 We take some vector $(y_{0},y_{1},\ldots y_{N},y_{N+1})$, multiply the equation in (1.4) by $y_{n}$, sum up and change the order of summation:
@@
   • 3)
-    Calculate parameters $d_{i}$, $\rho_{i}$, $H_{ki}$, $\omega_{i}$ using factorization (2.2) and formulas (2.7), (2.8), (2.12).
+    Calculate parameters $d_{i}$, $\rho_{i}$, $H_{ki}$, $\omega_{i}$ using factorization (2.2) and formulas (2.4), (2.8), (2.9), (2.13).
   • 4)